OpenStudy (anonymous):

Factor by grouping: mz-5mh^2-5nz+25nh^2 "@whpalmer4"

OpenStudy (anonymous):

I meant to add in factor by grouping

OpenStudy (anonymous):

"@whpalmer4

OpenStudy (whpalmer4):

\[mz-5mh^2-5nz+25nh^2\] \[(mz-5mh^2) -(5nz - 25nh^2)\](notice that the negative sign appears in front of \(25nh^2\) because I pulled \(-1\) out as a factor) \[m(z-5h^2) - 5n(z-5h^2)\]Common factor here is \((z-5h^2)\) so factor again:\[(z-5h^2)(m-5n)\] Test our work: \[(z-5h^2)(m-5n) = z*m -5n*z -5h^2*m - 5h^2(-5n)\]\[=mz-5nz-5h^2m+25h^2n = mz-5mh^2-5nz+25nh^2\checkmark\]

OpenStudy (anonymous):

ok thanks so much